Determine the equation of each line. Write the answer in slope intercept form.
4. (52) The line through (4, 0) that is perpendicular to the line x + y = 3
I can't figure out where to start graphing? Why can't I understand this?

Answer by stanbon(75887)

The given line has a slpe of -1
So the line perpendicular to it has a slope of 1
The equation you want has the form y=mx+b
You are told x=4, y=0, m=1
Substituting you can find "b", as follows:
0=1(4)+b
b=-4
The equation you want is y=x-4
Cheers,
Stan H.